(2+3i)(6+4i)=12+8i+18i+12i^2
=12+26i-12 (since i^2 = -1)
=26i ans

Answer by vksarvepalli(154)

(2+3i)(6+4i)=12+8i+18i+12i^2
=12+26i-12 (since i^2 = -1)
=26i ans